OpenStudy (anonymous):

Which system of equations is represented by the graph? y=-x-10/x-4 y=x

OpenStudy (anonymous):

y=-x-10/x-4, y=x y=x+10/x-4, y=-x y=-x+10/x-4, y=x y=-x+10/x-4, y=-x

OpenStudy (anonymous):

Your answer would be the first one :]!

OpenStudy (anonymous):

Where I don't see it

OpenStudy (anonymous):

@enchanted_bubbles explain it ta him

OpenStudy (anonymous):

The answer is your first option.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Your welcome :]